A tower, 28.4ft high must be secured with a guy wire anchored 5ft. from the base of the

tower. What angle will the guy wire make with the ground?


1
Expert's answer
2022-01-17T16:04:26-0500

let"\theta" be the angle between the wire and the ground.

"tan\theta =opp/adj"

"tan\theta" =28.4/5

"tan\theta=5.68"

"tan^-1 (tan\theta)=tan^-1(5.68)"

"\theta" =80.02